OpenStudy (anonymous):

10. triangle ABC has side lengths 8, 15, and 17. Do the side lengths form a Pythagorean triple? (1 point) yes no ?

OpenStudy (raden):

is 17^2 = 8^2 + 15^2 ?

OpenStudy (anonymous):

?

OpenStudy (raden):

you just check it, if 17^2 = 8^2 + 15^2 then it is the Pythagorean triple

OpenStudy (anonymous):

thx

OpenStudy (raden):

welcome

OpenStudy (anonymous):

Its Yes